OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = one over fourx + 7 and that pass through the point (−8, 9). parallel: y = one over fourx − 11 perpendicular: y = −4x + 23 parallel: y = −4x + 11 perpendicular: y = one over fourx − 4 parallel: y = 4x − 7 perpendicular: y = negative one over fourx − 7 parallel: y = one over fourx + 11 perpendicular: y = −4x − 23 @mathstudent55

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Parallel lines have the same slope. When you write the equation of a line as y = mx + b, m is the slope. Perpendicular lines have slopes that are negative reciprocals. That means if you know the slope of a line, the slope of a perpendicular is obtained by flipping the slope and changing the sign.

OpenStudy (anonymous):

ok.....

OpenStudy (mathstudent55):

Your line is \(y = \dfrac{1}{4}x + 7 \) The slope is \(m = \dfrac{1}{4}\) A parallel line has slope \( m = \dfrac{1}{4} \), the same. A perpendicular line has slope \(m = -4\), the negative reciprocal. Only choices A and D have the correct slopes for the lines, so eliminate B and C.

OpenStudy (anonymous):

ok I think its B? am I right/

OpenStudy (mathstudent55):

? Do you read what I write?

OpenStudy (anonymous):

i meant A

OpenStudy (mathstudent55):

Let's test A. \(y = \dfrac{1}{4}x - 11 \) Test point (-8, 9) \(9 = \dfrac{1}{4}(8) - 11 \) \( 9 = 2 - 11 \) 9 = -2 Not true. This can't be the correct choice, so the correct choice must be D.